OpenStudy (anonymous):

The denominator of a fraction is 2 more then the numerator. If both denominator and numerator are increased by 1 the fraction becomes 2/3. Find the original fraction

OpenStudy (lgbasallote):

let x = numerator \(\LARGE \frac{x}{x+2}\) is the fraction \(\LARGE \frac{x+1}{x+3} = \frac{3}{4}\) cross multiply... 4x + 4 = 3x + 9 find x

Parth (parthkohli):

\(\Large \color{purple}{{x + 1 \over x + 3} = {3 \over 4}} \)

OpenStudy (lgbasallote):

do you knoow the answer @Hollywood_chrissy ? you can verify with me :)

Parth (parthkohli):

\(\Large \color{purple}{ 4x = 3x + 5 } \) \(\Large \color{purple}{ 4x - 3x = 5 } \) \(\Large \color{purple}{ x = 5 } \)

OpenStudy (anonymous):

do not know if you noticed but I actually make an error in the question, that fraction is 2/3 nor 3/4 as I first wrote.

Parth (parthkohli):

I see how you have edited the question: \(\Large \color{purple}{ {x +1 \over x + 3} = {2 \over 3} } \)

Parth (parthkohli):

\(\Large \color{purple}{ 2x + 6 = 3x + 3 } \)

OpenStudy (lgbasallote):

^that

Parth (parthkohli):

\(\Large \color{purple}{ -x = -3 } \) \(\Large \color{purple}{ x = ? } \)

OpenStudy (anonymous):

3/5 ?

OpenStudy (anonymous):

I got 3/5

Parth (parthkohli):

WHUT? Oh God.

Parth (parthkohli):

I yeah lol

OpenStudy (lgbasallote):

and 3/5 seems to find the condition though

Parth (parthkohli):

It's actually: \(\Large \color{purple}{ -x = -3 } \) \(\Large \color{purple}{ x = 3 } \)

OpenStudy (lgbasallote):

oh oh yeah...that makes more sense \(\huge \frac{3}{5} \leftarrow \text{final answer}\)